9-55 Year E Sens 0 $90 -$110-$100 $120 35 35 35 35 9-60 20 2 20 20 4 20 5 20 6 20 10 20 30 40 50 180

9-55 Year E Sens 0 $90 -$110-$100 $120 35 35 35 35 9-60 20 2 20 20 4 20 5 20 6 20 10 20 30 40 50 180
)Based on the payback period, which altermnative-41 Amal (b) Based on future worth analysis, which of the (c) Based on future worth analysis, which alterna- (a) At 10% interest, what is the benefit-cost ratio is preferred? four alternatives is preferred at 5% interest? tive is preferred at 20% interest? for Alt. G? change (a) W (b) V 9-62 Co tory equipment costing
Answer #1

a) Based on PBP, alt F is a better option

Year

E

Cumulative Cash flow

F

Cumulative Cash flow

G

Cumulative Cash flow

H

Cumulative Cash flow

0

-90

-90

-110

-110

-100

-100

-120

-120

1

20

-70

35

-75

0

-100

0

-120

2

20

-50

35

-40

10

-90

0

-120

3

20

-30

35

-5

20

-70

0

-120

4

20

-10

35

30

30

-40

0

-120

5

20

10

0

30

40

0

0

-120

6

20

30

0

30

50

50

180

60

PBP

4.50

3.14

5.00

6.00

The formula = negative last cumulative cash flow period+(abs(negative last cumulative cash flow/positive cash flow of next period)

So, here for alt F, PBP = 3+(5/35) = 3.14

b. Based on FW, alt H is a better option

FW- Compounded cash flow at 5%

Year

E

F

G

H

0

-90.00

-110.00

-100.00

-120.00

1

21.00

36.75

0.00

0.00

2

22.05

38.59

11.03

0.00

3

23.15

40.52

23.15

0.00

4

24.31

42.54

36.47

0.00

5

25.53

0.00

51.05

0.00

6

26.80

0.00

67.00

241.22

FW

52.84

48.40

88.70

121.22

The formula is cashflow*(1+0.05)^n, n is number of years

c Here too, alt H is a better option

FW- Compounded cash flow at 20%

Year

E

F

G

H

0

-90.00

-110.00

-100.00

-120.00

1

24.00

42.00

0.00

0.00

2

28.80

50.40

14.40

0.00

3

34.56

60.48

34.56

0.00

4

41.47

72.58

62.21

0.00

5

49.77

0.00

99.53

0.00

6

59.72

0.00

149.30

537.48

FW

148.32

115.46

260.00

417.48

The formula is cashflow*(1+0.2)^n, n is number of years

d. Alt G B:C ratio is 0.97

Year

G

Discounted Benefit

Discounted Cost

0

-100

100

1

0

0.00

2

10

8.26

3

20

15.03

4

30

20.49

5

40

24.84

6

50

28.22

96.84

100

B:C = Sum of discounted Benefit/ Sum of discounted Cost = 96.84/100 = 0.9684

Discounting formula = Cashflow/(1+0.1)^n , n is number of years